OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 9 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth. I need help with this to be explained

OpenStudy (imstuck):

39.9Ok, here's the formula for the volume of a cone\[V=\frac{ 1 }{ 3 }\pi r ^{2}h\] You are given the volume as 300, you are given the radius as 9, so just fill in the formula using that to get \[300=\frac{ 1 }{ 3 }\pi (9)^{2}h\] Now solve for h. Multiply both sides by 3 to get rid of the denominator in the one-third. Now you have 900 = pi (81)h. Divide both sides by 81p and you get\[\frac{ 900 }{ 81\pi }=h\] Dividing this out you get the height as 11.1pi. You now have to multiply the 11.11 ties pi to get 39.906 or rounded to 39.9

OpenStudy (anonymous):

Thank you! I get the problem now!

OpenStudy (imstuck):

Yay!!!!!